OpenStudy (anonymous):

find s for the geometric series 4+12+36+108+. 20

OpenStudy (zale101):

first, you must use the geometric series formula for Sn \[Sn= \frac{ a(r^n - 1) }{ (r-1) }\]

OpenStudy (zale101):

since S(20), that means n = 20 replace all the n in the formula to 20 and since the series is 3 times the previous number 4*3 =12 12*3=36 36*3=108 then r=3 and a=4 4 sequence

OpenStudy (zale101):

\[S_(20)=\frac{ 4(3^{20}-1 )}{ 3-1}\]